OpenStudy (anonymous):

If f(x) = x2 - 4x + 1, how would i find the range of f(x) ??

OpenStudy (anonymous):

find the second coordinate of the vertex and then say from there to infinity

OpenStudy (anonymous):

x^2 is a parabola. Find the min value or vertex. From min to...(you know how a parabola behaves)

OpenStudy (anonymous):

magic formula for vertex is \[-\frac{b}{2a}=-\frac{-4}{2}=2\] \[f(2)=2^2-4\times 2+1=-3\] so range is \[[-3,\infty)\]

OpenStudy (anonymous):

\[2x ^{2}+5x-3\le0\] I have another one how do i go about solving this?

OpenStudy (anonymous):

find the zeros. it will be negative between them

OpenStudy (anonymous):

\[2x^2+5x-3=0\] \[(2x-1)(x+3)=0\] \[x=-3\] or \[x=\frac{1}{2}\]

OpenStudy (anonymous):

so negative on the interval \[(-3,\frac{1}{2})\]